# Long-term Causal Inference via Modeling Sequential Latent Confounding

### v1

@@ Abstract
Long-term causal inference is an important but challenging problem across various scientific domains. To solve the latent confounding problem in long-term observational studies, existing methods leverage short-term experimental data. Ghassami et al. (2022) propose an approach based on the Conditional Additive Equi-Confounding Bias (CAECB) assumption, which asserts that the confounding bias in the short-term outcome is equal to that in the long-term outcome, so that the long-term confounding bias and the causal effects can be identified. While effective in certain cases, this assumption is limited to scenarios with a one-dimensional short-term outcome. In this paper, we introduce a novel assumption that extends the CAECB assumption to accommodate temporal short-term outcomes. Our proposed assumption states a functional relationship between sequential confounding biases across temporal short-term outcomes, under which we theoretically establish the identification of long-term causal effects. Based on the identification result, we develop an estimator and conduct a theoretical analysis of its asymptotic properties. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
Long-term causal inference is an important but challenging problem across various scientific fields, such as education (Athey et al., 2019), medicine (Fleming et al., 1994), and marketing (Hohnhold et al., 2015). While in many real-world scenarios, long-term observational data are readily available, a major challenge for long-term causal inference is the presence of latent confounding in observational studies. A common way to mitigate this issue is incorporating short-term experimental data, which raises a fundamental question: how can short-term experimental data be leveraged to address latent confounding in observational data for long-term causal inference, as shown in Figure 1?
Figure 1: Causal graphs of experimental data and observational data with $\mathbf{X}$ being covariates, $\mathbf{U}$ being latent confounders, $A$ being treatment, $\mathbf{S}$ being short-term outcome, and $Y$ being long-term outcome. Fig. 1(a) represents the causal graph of the short-term experimental data, where treatment $A$ is not affected by latent confounders $\mathbf{U}$ and the long-term outcome $Y$ is unobserved. Fig. 1(b) represents the causal graph of the long-term observational data, where the latent confounders $\mathbf{U}$ affect treatments $A$ and outcomes $\mathbf{S},Y$ and the long-term outcome $Y$ can be observed. Fig. 1(c) represents the overall graph with expanding temporal short-term outcomes.
@@
Existing works explore various methods to mitigate latent confounding by utilizing observational and experimental data based on different assumptions. One widely used assumption is the Latent Unconfoundedness assumption (Athey et al., 2020; Chen & Ritzwoller, 2023; Yang et al., 2024), which posits the short-term potential outcomes mediate the long-term potential outcome in the observational data, i.e., $Y(a){\perp\!\!\!\perp}A|\mathbf{S}(a),\mathbf{X},G=O$. These methods are effective when short-term outcomes $\mathbf{S}$ contain substantial information about the latent confounder $\mathbf{U}$, such as in studies on the lifetime effects of youth employment and training programs in the United States (Aizer et al., 2024). However, this assumption essentially restricts that the latent confounders can not affect the long-term outcome, i.e., ruling out the causal edge $\mathbf{U}\rightarrow Y$. To address this limitation, a follow-up work (Ghassami et al., 2022) proposes a novel assumption called Conditional Additive Equi-Confounding Bias (CAECB), allowing for the causal link $\mathbf{U}\rightarrow Y$. The CAECB assumption states that the confounding bias in the short-term outcome is equal to that in the long-term outcome, enabling the identification of long-term causal effects. This assumption may be more reasonable and more aligned with practical settings where confounding biases share a similar manner over short- and long-term outcomes conditional on covariates $\mathbf{X}$, such as in research studying the impact of school finance reforms on student outcomes (Jackson, 2016).
@@
Following the work of Ghassami et al. (2022), we focus on the setting where the causal link $\mathbf{U}\rightarrow Y$ exists. Beyond this, we further consider temporal short-term outcomes, as shown in Figure 1(c), a more common scenario where the method under the CAECB assumption (Ghassami et al., 2022) is not applicable. In many real-world applications, short-term outcomes exhibit temporal dependencies, and capturing these relationships is essential for inferring long-term causal effects. For example, in evaluating the long-term (e.g. year-long) effectiveness of the medication, patients undergo regular follow-up visits (e.g., weekly or monthly). During these visits, temporal short-term health indicators are recorded as short-term outcomes. These sequential measurements capture the progression of patients’ conditions over time, providing valuable information on the long-term outcome of interest. Consequently, the absence of temporal considerations in existing methods limits their effectiveness and constrains the potential of long-term causal inference in practice.
@@ 2 Related Work
Long-term causal inference For years, researchers have investigated which short-term outcomes can reliably predict long-term causal effects. Various criteria have been proposed for identifying valid surrogates, including the Prentice criteria (Prentice, 1989), principal surrogacy (Frangakis & Rubin, 2002), strong surrogate criteria (Lauritzen et al., 2004), causal effect predictiveness (Gilbert & Hudgens, 2008), and consistent surrogate and its variations (Chen et al., 2007; Ju & Geng, 2010; Yin et al., 2020). Recently, there has been growing interest in estimating long-term causal effects using surrogates, which is also the focus of this paper. One line of work assumes the unconfoundedness assumption. Under the unconfoundedness assumption, LTEE (Cheng et al., 2021) and Laser (Cai et al., 2024) are based on specifically designed neural networks for long-term causal inference. EETE (Kallus & Mao, 2020) explores the data efficiency from the surrogate in several settings and proposes an efficient estimator for treatment effect. ORL (Tran et al., 2023) introduces a doubly robust estimator for average treatment effects using only short-term experiments, additionally assuming stationarity conditions between short-term and long-term outcomes. Singh & Sun (2024) proposes a kernel ridge regression-based estimator for long-term effect under continuous treatment. Additionally, Wu et al. (2024) develop a policy learning method for balancing short-term and long-term rewards. Our work is different from them. We do not assume the unconfoundedness assumption, and we use the data combination technique to solve the problem of unobserved confounders. Another line of research, which avoids the unconfoundedness assumption, tackles the latent confounding problem by combining experimental and observational data. This setting is first introduced by the method proposed by Athey et al. (2019), which, under surrogacy assumption, constructs the so-called Surrogate Index (SInd) as the substitutions for long-term outcomes in the experimental data for effect identification. As follow-up work, Athey et al. (2020) introduces the latent unconfoundedness assumption, which assumes that short-term potential outcomes can mediate the long-term potential outcomes, thereby enabling long-term causal effect identification. Under this assumption, Yang et al. (2024); Chen & Ritzwoller (2023) propose several estimators for effect estimation. The alternative feasible assumptions (Ghassami et al., 2022) are proposed to replace the latent unconfoundedness assumption, e.g., the Conditional Additive Equi-Confounding Bias (CAECB) assumption. Based on proximal methods (Tchetgen Tchetgen et al., 2024), Imbens et al. (2024) proposes considering the short-term outcomes as proxies of latent confounders, thereby achieving effect identification. However, these studies primarily concentrate on average treatment effects, whereas our focus in this paper is on heterogeneous effects. Among the existing literature, the most closely related work is Ghassami et al. (2022)’s work. Our work can be viewed as a significant extension of theirs, as we consider a more practical scenario where short-term outcomes exhibit temporal dependencies, and theoretically, their CAECB assumption is a special case of the assumption we propose.
@@
Modeling Latent Confounding Bias An effective way to solve the latent confounding problem in the data combination setting is to model latent confounding bias. Kallus et al. (2018) proposes modeling confounding bias under a linearity assumption. Hatt et al. (2022) introduce to model the nonlinear confounding bias using the representation learning technique. Wu & Yang (2022) propose the integrative R-learner via a regularization for the conditional effects and confounding bias with the Neyman orthogonality. Zhou et al. (2025) propose a two-stage representation learning strategy to model such a confounding bias. Different from these works, we focus on the long-term causal inference setting, and rather than focus on how to model the confounding bias, we concentrate more on the relationship between sequential confounding biases.
@@ 3 Preliminaries
3.1 Notations, Problem Definition, Assumptions
Let $A\in\{0,1\}$ be the treatment variable, $\mathbf{X}\in{\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ be the observed covariates where $d$ is the dimension of $\mathbf{X}$, $\mathbf{U}\in{\mathcal{U}}\subseteq{\mathbb{R}}^{d_{U}}$ be the latent confounders where $d_{U}$ is the dimension of $\mathbf{U}$. Let $\mathbf{S}=[S_{1},S_{2},\dots,S_{T}]$ be the short-term outcome variable where $S_{t}\in{\mathcal{S}}\subseteq{\mathbb{R}}$ is the short-term outcome measured at timestep $t$, and $Y=S_{T+\mu}\in{\mathcal{R}}\subseteq{\mathbb{R}}$ be the long-term outcome. Further, we leverage the potential outcome framework proposed by Rubin (1978). We denote $\mathbf{S}(a)$ as the potential short-term outcome, $S_{t}(a)$ as the potential short-term outcome at timestep $t$, and $Y(a)$ as the potential long-term outcome. Following existing work on long-term inference (Athey et al., 2020; Ghassami et al., 2022; Imbens et al., 2024; Chen & Ritzwoller, 2023), we denote $G\in\{E,O\}$ as the indicator of data group, where $G=E$ indicates experimental data and $G=O$ indicates observational data. Let lowercase letters (e.g., $a,\mathbf{x},\mathbf{u},\mathbf{s},y,s(a),y(a)$) denote the value of the previously described random variables. Let the superscript $(i)$ denote a specific unit, e.g., $\mathbf{x}^{(i)}$ is the covariate value of unit $i$. Then, the experimental data and the observational data are denoted as $\mathbb{D}_{e}=\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},g^{(i)}=E\}_{i=1}^{n_{e}}$ and $\mathbb{D}_{o}=\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},y^{(i)},G^{(i)}=O\}_{i=n_{e}+1}^{n_{e}+n_{o}}$, where $n_{e}$ and $n_{o}$ are the size of experimental data and the observational data respectively.
@@ 3.1 Notations, Problem Definition, Assumptions
$\mu_{S_{t}}^{E}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}|A,\mathbf{X},G=E],$ (1)
$\mu_{S_{t}}^{O}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}|A,\mathbf{X},G=O],$
$\mu_{Y}^{E}(A,\mathbf{X})={\mathbb{E}}[Y|A,\mathbf{X},G=E],$
$\mu_{Y}^{O}(A,\mathbf{X})={\mathbb{E}}[Y|A,\mathbf{X},G=O],$
@@
where $\omega_{t}(\mathbf{X})$ is known as the confounding bias, the discrepancy between the conditional mean outcome differences derived from the experimental data and the observational data (see Sections 3.2 and 4 for more details on how it serves to the identification of long-term effects).
@@
Task: Given a short-term experimental dataset $\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},g^{(i)}=E\}_{i=1}^{n_{e}}$ and a long-term observational dataset $\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},y^{(i)},g^{(i)}=O\}_{i=n_{e}+1}^{n_{e}+n_{o}}$, the goal is to identify the following causal estimand of interest: the Heterogeneous Long-term Causal Effects (HLCE), i.e.,
$\tau(\mathbf{x})={\mathbb{E}}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}].$ (2)
However, HLCE $\tau(\mathbf{x})$ is not identifiable from the experimental data alone, since $Y$ is missing in that dataset. Also it is not identifiable from the observational data alone, since the observational data suffers from the latent confounding problem. Furthermore, the information regarding the causal effects in the experimental data is not necessarily relevant to that in the observational data without further assumptions. To ensure the identification, we first make the following assumptions that are commonly assumed in long-term causal inference:
@@
If a unit is assigned treatment, we observe its associated potential outcome. Formally, if $A=a$, then $Y=Y(a)$, and $\mathbf{S}=\mathbf{S}(a)$.
@@
The treatment assignment is non-deterministic. Formally, $\forall a,\mathbf{x}$, we have $10<P(A=a|\mathbf{X}=\mathbf{x})<$ 1, and $0<P(G=O|A=a,\mathbf{X}=\mathbf{x})<1$.
Assumption 3.3 (Weak internal validity of observational data).
Latent confounders exist in observational data. Formally, afor-$\forall a$, we have $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},\mathbf{U},G=O$ and $A\not{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},G=O$.
Assumption 3.4 (Internal validity of experimental data).
There are no latent confounders in experimental data. Formally, afor-$\forall a$, we have $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X},G=E$.
Assumption 3.5 (External validity of experimental data).
The distribution of the potential outcomes is invariant to whether the data belongs to the experimental or observational data. Formally, afor-$\forall a$, we have $G{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}|\mathbf{X}$.
Assumptions 3.1 and 3.2 are standard assumptions in causal inference (Rubin, 1974; Imbens, 2000). Assumptions 3.3, 3.4 and 3.5 are mild and widely assumed in data combination settings (Shi et al., 2023; Imbens et al., 2024; Athey et al., 2019, 2020; Hu et al., ). Specifically, Assumption 3.3 allows the existence of latent confounders in observational data, thus it is much weaker than the traditional unconfoundedness assumption. Assumption 3.4 is reasonable and can hold since the treatment assignment mechanism is under control in the experiments. Assumption 3.5 connects the potential outcome distributions between observational and experimental data.
Assumptions 3.1, 3.2, 3.3, 3.4 and 3.5 are still not sufficient to identify the causal estimand of interest. The root cause is that, even though the assumptions above link the experimental and observational data, the (long-term) latent confounding problem remains unsolved. In the following section, we first review a method proposed by Ghassami et al. (2022), which poses an extra assumption, called Conditional Additive Equi-Confounding Bias (CAECB) assumption, to achieve the identification of HLCE. Then, we propose our approach that generalizes the method of Ghassami et al. (2022) to allow temporal short-term outcomes.
(a) Schematic representation of CAECB assumption (Ghassami et al., 2022).
@@
Figure 2: Schematic representations of CAECB assumption proposed by Ghassami et al. (2022) and our FCAECB assumption. As shown in Figure 2(a), the CAECB assumption requires that the confounding bias in the short-term outcome is equal to that in the long-term outcome. As shown in Figure 2(b), the FCAECB assumption relaxes this constraint by allowing for temporal short-term outcomes and only requiring that confounding biases across different time steps follow a specific pattern rather than remaining equal.
3.2 CAECB Assumption (Ghassami et al., 2022)
Ghassami et al. (2022) introduced a method for combining experimental and observational data to identify the HLCE under an extra CAECB assumption. The method under CAECB assumption is only applicable to scenarios with a one-dimensional short-term outcome, which we denote as $S$ (and its corresponding potential outcome $S(a)$) with slightly abusing notation.
To begin with, we first restate the CAECB assumption as follows:
Assumption 3.6 (Conditional Additive Equi-Confounding Bias, CAECB (Ghassami et al., 2022)).

### v2

@@ Abstract
Long-term causal inference is an important but challenging problem across various domains. To solve the latent confounding problem in long-term observational studies, existing methods leverage short-term experimental data. Ghassami et al. [14] propose an approach based on the Conditional Additive Equi-Confounding Bias (CAECB) assumption, which asserts that the confounding bias in the short-term outcome is equal to that in the long-term outcome, so that the long-term confounding bias and the causal effects can be identified. While effective in certain cases, this assumption is limited to scenarios where there is only one short-term outcome with the same scale as the long-term outcome. In this paper, we introduce a novel assumption that extends the CAECB assumption to accommodate temporal short-term outcomes. Our proposed assumption states a functional relationship between sequential confounding biases across temporal short-term outcomes, under which we theoretically establish the identification of long-term causal effects. Based on the identification result, we develop an estimator and conduct a theoretical analysis of its asymptotic properties. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
Long-term causal inference is an important but challenging problem across various scientific fields, such as education [4], medicine [11], and marketing [17]. While in many real-world scenarios, long-term observational data are readily available, a major challenge for long-term causal inference is the presence of latent confounding in observational studies. A common way to mitigate this issue is to incorporate short-term experimental data, which raises a fundamental question: how can short-term experimental data be leveraged to address latent confounding in observational data for long-term causal inference, as shown in Fig. 1(a) and 1(b)?
@@
Existing works explore various methods to mitigate latent confounding by utilizing observational and experimental data based on different assumptions  [^1] Please kindly refer to Appendix A for detailed discussion on more existing assumptions.. One widely used assumption is the Latent Unconfoundedness assumption [5, 8, 28], which posits the short-term potential outcomes mediate the long-term potential outcome in the observational data, i.e., $Y(a){\perp\!\!\!\perp}A\mid\mathbf{S}(a),\mathbf{X},G=O$. These methods are effective when short-term outcomes $\mathbf{S}$ contain substantial information about the latent confounder $\mathbf{U}$, such as in studies on the lifetime effects of youth employment and training programs in the United States [1]. However, this assumption essentially restricts that the latent confounders can not affect the long-term outcome, i.e., ruling out the causal edge $\mathbf{U}\rightarrow Y$. To address this limitation, a follow-up work [14] proposes a novel assumption called Conditional Additive Equi-Confounding Bias (CAECB), allowing for the causal link $\mathbf{U}\rightarrow Y$. The CAECB assumption states that the confounding bias in the short-term outcome is equal to that in the long-term outcome, enabling the identification of long-term causal effects. This assumption may be more reasonable and more aligned with practical settings where confounding biases share a similar manner over short- and long-term outcomes conditional on covariates $\mathbf{X}$, such as in research studying the impact of school finance reforms on student outcomes [21].
@@
Following the work of Ghassami et al. [14], we focus on the setting where the causal link $\mathbf{U}\rightarrow Y$ exists. Beyond this, we further consider temporal short-term outcomes, as shown in Figure 1(c), a more common scenario where the method under the CAECB assumption [14] is not applicable. In many real-world applications, short-term outcomes exhibit temporal dependencies, and capturing these relationships is essential for inferring long-term causal effects. For example, in evaluating the long-term (e.g. year-long) effectiveness of the medication, patients undergo regular follow-up visits (e.g., weekly or monthly). During these visits, temporal short-term health indicators are recorded as short-term outcomes. These sequential measurements capture the progression of patients’ conditions over time, providing valuable information on the long-term outcome of interest. Consequently, the absence of temporal considerations in existing methods limits their effectiveness and constrains the potential of long-term causal inference in practice.
Figure 1: Causal graphs for experimental and observational data with $\mathbf{X}$ as covariates, $\mathbf{U}$ as latent confounders, $A$ as treatment, $\mathbf{S}$ as short-term outcomes, and $Y$ as the long-term outcome. Fig. 1(a) shows the experimental data, where $A$ is unaffected by $\mathbf{U}$ and $Y$ cannot be unobserved. Fig. 1(b) depicts the observational data, where $\mathbf{U}$ affects $A$, $\mathbf{S}$, and $Y$, and $Y$ can be observed. Fig. 1(c) illustrates the full causal graph with temporally extended short-term outcomes.
@@ 2 Related Work
Long-term causal inference For years, researchers have investigated which short-term outcomes can reliably predict long-term causal effects [30, 12, 27, 15, 7, 22, 42]. Recently, there has been growing interest in estimating long-term causal effects using surrogates, which is also the focus of this paper. One line of work assumes the unconfoundedness assumption. Under the unconfoundedness assumption, LTEE [9] and Laser [6] are based on specifically designed neural networks for long-term causal inference. EETE [23] explores the data efficiency from the surrogate in several settings and proposes an efficient estimator for treatment effects. ORL [37] introduces a doubly robust estimator for average treatment effects using only short-term experiments, additionally assuming stationarity conditions between short-term and long-term outcomes. Singh and Sun [34] proposes a kernel ridge regression-based estimator for long-term effect under continuous treatment. Additionally, Yang et al. [40], Wu et al. [39] develop a policy learning method for balancing short-term and long-term rewards. Our work is different from theirs. We do not rely on the unconfoundedness assumption, and we use the data combination technique to solve the problem of unobserved confounders. Another line of work, which avoids the unconfoundedness assumption, tackles the latent confounding problem by combining experimental and observational data. This setting is first introduced by Athey et al. [4], which, under surrogacy assumption, constructs the so-called Surrogate Index (SInd) as the substitutions for long-term outcomes in the experimental data for effect identification. As follow-up work, Athey et al. [5] introduce the latent unconfoundedness assumption, which assumes that short-term potential outcomes can mediate the long-term potential outcomes, thereby enabling long-term causal effect identification. Under this assumption, Meza and Singh [28], Yang et al. [41], Chen and Ritzwoller [8] propose several estimators for effect estimation. The alternative feasible assumptions [14] are proposed to replace the latent unconfoundedness assumption, e.g., the Conditional Additive Equi-Confounding Bias (CAECB) assumption. Based on proximal methods [36], Imbens et al. [19] propose considering the short-term outcomes as proxies of latent confounders, thereby achieving effect identification. Among the existing literature, the most closely related work is the work of Ghassami et al. [14]. Our work can be viewed as a significant extension of theirs, as we consider a more practical scenario where short-term outcomes exhibit temporal dependencies, and theoretically, their CAECB assumption is a special case of our proposed assumption. Our work is also closely related to the work of Imbens et al. [19], which also explores the sequential short-term outcomes and assumes them as proxies of latent confounders. Different from that, we explore the sequential information from short-term confounding bias, which allows for inferring long-term confounding bias and results in the final effect identification.
@@
Modeling Latent Confounding Bias An effective way to solve the latent confounding problem in the data combination setting is to model latent confounding bias. Kallus et al. [24] proposes modeling confounding bias under a linearity assumption. Hatt et al. [16] introduce to model the nonlinear confounding bias using the representation learning technique. Wu and Yang [38] propose the integrative R-learner via a regularization for the conditional effects and confounding bias with the Neyman orthogonality. Zhou et al. [43] propose a two-stage representation learning strategy to model such a confounding bias. Different from these works, we focus on the long-term causal inference setting, and rather than focus on how to model the confounding bias, we concentrate more on the relationship between sequential confounding biases.
@@ 3 Preliminaries
3.1 Notations, Problem Definition, Basic Assumptions
Let $A\in\{0,1\}$ be the treatment variable, $\mathbf{X}\in{\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ be the observed covariates where $d$ is the dimension of $\mathbf{X}$, $\mathbf{U}\in{\mathcal{U}}\subseteq{\mathbb{R}}^{d_{U}}$ be the latent confounders where $d_{U}$ is the dimension of $\mathbf{U}$. Let $\mathbf{S}=[S_{1},S_{2},\dots,S_{T}]$ be the short-term outcome variable where $S_{t}\in{\mathcal{S}}\subseteq{\mathbb{R}}$ is the short-term outcome measured at time step $t$, and $Y=S_{T+\mu}\in{\mathcal{R}}\subseteq{\mathbb{R}}$ be the long-term outcome. Further, we leverage the potential outcome framework [32]. We denote $\mathbf{S}(a)$ as the potential short-term outcome, $S_{t}(a)$ as the potential short-term outcome at time step $t$, and $Y(a)$ as the potential long-term outcome. Following existing work on long-term inference [5, 14, 19, 8], we denote $G\in\{E,O\}$ as the indicator of data group, where $G=E$ indicates experimental data and $G=O$ indicates observational data. Let lowercase letters (e.g., $a,\mathbf{x},\mathbf{u},\mathbf{s},y,s(a),y(a)$) denote the value of the previously described random variables. Let the superscript $(i)$ denote a specific unit, e.g., $\mathbf{x}^{(i)}$ is the covariate value of unit $i$. Then, the experimental data and the observational data are denoted as $\mathbb{D}_{e}=\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},g^{(i)}=E\}_{i=1}^{n_{e}}$ and $\mathbb{D}_{o}=\{a^{(i)},\mathbf{x}^{(i)},\mathbf{s}^{(i)},y^{(i)},G^{(i)}=O\}_{i=n_{e}+1}^{n_{e}+n_{o}}$, where $n_{e}$ and $n_{o}$ are the size of experimental data and the observational data respectively.
@@ 3.1 Notations, Problem Definition, Basic Assumptions
$\mu_{S_{t}}^{E}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}\mid A,\mathbf{X},G=E],\quad\mu_{S_{t}}^{O}(A,\mathbf{X})={\mathbb{E}}[\mathbf{S}_{t}\mid A,\mathbf{X},G=O],$ (1)
$\mu_{Y}^{E}(A,\mathbf{X})={\mathbb{E}}[Y\mid A,\mathbf{X},G=E],\quad\mu_{Y}^{O}(A,\mathbf{X})={\mathbb{E}}[Y\mid A,\mathbf{X},G=O],$
@@
where $\omega_{t}(\mathbf{X})$ is known as the confounding bias, the discrepancy between the conditional mean outcome differences derived from the experimental data and the observational data (Please see Sections 3.2 and 4 for more details on how it serves to the identification of long-term effects).
@@
Task: Given a short-term experimental dataset $\mathbb{D}_{e}$ and a long-term observational dataset $\mathbb{D}_{o}$, the goal in this paper is to identify and estimate the Heterogeneous Long-term Causal Effects (HLCE), i.e.,
$\tau(\mathbf{x})={\mathbb{E}}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}].$ (2)
However, HLCE $\tau(\mathbf{x})$ is not identifiable from the experimental data alone, since $Y$ is missing in that dataset. Also it is not identifiable from the observational data alone, since the observational data suffers from the latent confounding problem, i.e., $A\not{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}\mid\mathbf{X},G=O$. Furthermore, the information regarding the causal effects in the experimental data is not necessarily relevant to that in the observational data without further assumptions. To ensure the identification, we first make the following assumptions that are commonly assumed in long-term causal inference [4, 5, 8, 14, 19]:
@@
If $A=a$, then $Y=Y(a)$, and $\mathbf{S}=\mathbf{S}(a)$.
@@
$\forall a,\mathbf{x}$, $0<P(A=a\mid\mathbf{X}=\mathbf{x});P(G=O\mid A=a,\mathbf{X}=\mathbf{x})<1$.
Assumption 3.3 (Observational Data).
afor-$\forall a$, $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}\mid\mathbf{X},\mathbf{U},G=O$.
Assumption 3.4 (Experimental Data).
afor-$\forall a$, $A{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}\mid\mathbf{X},G=E$.
Assumption 3.5 (Data Combination).
afor-$\forall a$, $G{\perp\!\!\!\perp}\{Y(a),\mathbf{S}(a)\}\mid\mathbf{X}$.
Assumptions 3.1 and 3.2 are standard assumptions in causal inference [31, 20]. Assumptions 3.3, 3.4 and 3.5 are mild and widely assumed in data combination settings [33, 19, 4, 5, 18]. Specifically, Assumption 3.3 allows the existence of latent confounders in observational data, thus it is much weaker than the traditional unconfoundedness assumption. Assumption 3.4 is reasonable and can hold since the treatment assignment mechanism is under control in the experiments. Assumption 3.5 connects the potential outcome distributions between observational and experimental data.
However, Assumptions 3.1, 3.2, 3.3, 3.4 and 3.5 are still not sufficient to identify the causal estimand of interest. The root cause is that, even though the assumptions above link the experimental and observational data, the (long-term) latent confounding problem remains unsolved. In the following section, we first review a method proposed by Ghassami et al. [14], which poses an extra assumption, called the Conditional Additive Equi-Confounding Bias (CAECB) assumption, to achieve the identification of HLCE. Then, we propose our approach that generalizes the method of Ghassami et al. [14] to allow temporal short-term outcomes.
(a) Schematic representation of CAECB assumption [14].
@@
Figure 2: Schematic representations of CAECB assumption proposed by Ghassami et al. [14] and our FCAECB assumption. As shown in Figure 2(a), the CAECB assumption requires that the confounding bias in the short-term outcome is equal to that in the long-term outcome. As shown in Figure 2(b), the FCAECB assumption relaxes this constraint by allowing for temporal short-term outcomes and only requiring that confounding biases across different time steps follow a specific pattern rather than remaining equal.
3.2 CAECB Assumption [14]
Ghassami et al. [14] introduced a method for combining experimental and observational data to identify the HLCE under an extra CAECB assumption. The method under CAECB assumption is only applicable to scenarios with a one-dimensional short-term outcome, which we denote as $S$ (and its corresponding potential outcome $S(a)$) with slightly abusing notation.
To begin with, we first restate the CAECB assumption:
Assumption 3.6 (Conditional Additive Equi-Confounding Bias, CAECB [14]).